Circular No. 8828

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 G1 (LINEAR) An apparently asteroidal object discovered by the LINEAR survey

(discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found by other astrometrists to have a cometary appearance. G. Sostero and E. Guido (Castellammare di Stabia, Italy, observing remotely with a 0.25-m f/3.4 reflector near Mayhill, NM) write that co-adding of 23 unfiltered 60-s exposures from Apr. 11.48 UT shows a small coma of total mag 17.8 nearly 10′′ in diameter, elongated toward p.a. 260^{o}; additional co-added 60-s images by Guido, A. Lepardo, and Sostero (0.45-m f/4.4 reflector at Remanzacco, Italy) on Apr. 11.91 show a faint coma nearly 8′′ in diameter. R. Ligustri (Udine, Italy, 0.32-m f/9 reflector) reports that ten 120-s images reveal a cometary appearance. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain reflector) reports that images taken on Apr. 12.05–12.15 at a scale of 1′′/pixel show a 5′′ coma, extended 6′′ on the southwest side (p.a. ∼ 240^{o}).

2007 UT  α_{2000}  δ_{2000}  Mag.
Apr. 10.37946  18^{h}05^{m}10.^{s}44  +33^{o}53′44′′.3  18.6

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-G37.

T = 2009 Mar. 8.637 TT  ω = 232.^{o}663
Ω = 72.236 }2000.0
q = 2.46334 AU  i = 94.461

SUPERNOVAE 2007at AND 2007au Two supernovae have been discovered on unfiltered CCD frames by W.

Li and N. Lee (LOSS/KAIT; cf. IAUC 8822). Discovery observations:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007at Mar. 18.53 14^{h}51^{m}12.^{s}28 −21^{o}23′31′′.2 17.7 7′′.7 E, 7′′.8 S
2007au Mar. 18.19 7 11 46.11 +49 51 13.4 17.0 41′′ E, 31′′ S

Additional information is available on the following Electronic Telegrams: 2007at in ESO 580-G44, CBETs 893, 898, 899, and 900 (type-Ia, discovered ∼ 9 days after maximum light); 2007au in UGC 3725, CBETs 895, 898, and 900 (type-Ia, discovered a couple of days before maximum).

2007 April 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green